OpenStudy (nerdybirdy6661):

ugh, i can't seem to get this adding and subtracting polynomials stuff. (7b^2−2b+3)−(5b^2−4+b^4 )−(9b^4+6) my answer was 10b^4+8b^2-2b+13

OpenStudy (evansanchez):

2b^2-10b^4-2b+1

OpenStudy (evansanchez):

answer

OpenStudy (nerdybirdy6661):

yeah i know the answer, i have the answer sheet, i just don't know how to get there

OpenStudy (lifeisadangerousgame):

Here's what you do first, nerdy. Distribute the subtraction signs to the polynomials so that you can get started. Keep everything as neat as possible. What I mean by that is multiply -1 to (5b^2−4+b^4 ) and (9b^4+6) and then write what you have

OpenStudy (lifeisadangerousgame):

After that, add/subtract like terms and see what you end up with. I have to go eat, I'll be back quite some time later :/

Directrix (directrix):

(7b^2−2b+3)−(5b^2−4+b^4 )−(9b^4+6) = 7b^2 - 2b + 3 - 5b^2 + 4 - b^4 - 9b^4 - 6 = -b^4 - 9b^4 +7b^2 - 5b^2 + 3 + 4 - 6 =

Directrix (directrix):

-b^4 - 9b^4 +7b^2 - 5b^2 + 3 + 4 - 6 = -10 b^4 + 2b^2 - 2b + 1